OpenStudy (anonymous):

Solve x2 + 9x + 9 = 0.

OpenStudy (anonymous):

hmm?

hero (hero):

You might have to either use complete the square or quadratic formula for this one

OpenStudy (anonymous):

can you show me how?

OpenStudy (anonymous):

x^2 + 9x +9=0 ?

jhonyy9 (jhonyy9):

x_1,_2=(-9+/- sqrt(81-36))/2 =(-9+/- sqrt45)/2 =(-9+/- 3sqrt5)/2 = = -3(3+/-sqrt5)/2